# Lower Limits on Scalar Ultralight Dark Matter from Optical Cavities

### v1

@@ Abstract
Oscillations of scalar ultralight dark matter (ULDM) at its Compton frequency would couple to fundamental constants to coherently drive the length of nearby Fabry-Perot cavities. Resulting differences in the length of two cavities are searched for in the spectrum of the beat note between lasers traversing the cavities. The new direct ULDM bounds set near 555\,5kHz and between 20 and 909090\,90kHz are one to two orders of magnitude lower for two model ULDM distributions – a standard galactic halo, and a relaxion star bound to Earth. Laser filtering and mechanical isolation are critical.
@@
This Letter reports new bounds on a well-motivated alternative – scalar, parity-even, ultralight dark matter (ULDM) that behaves as a classical wave [7]. A broadband search with the apparatus of Fig. 1 is for differing length changes of two cavities at the ULDM Compton frequency $f_{\phi}\equiv m_{\phi}c^{2}/(2\pi\hbar)$, for UDLM mass $m_{\phi}$, Planck constant $\hbar$, and speed of light $c$. Sensitivity to the local dark matter density produces strong and direct limits on boson stars [8, 9] composed of relaxions – a proposed solution to the hierarchy problem [10] – a contrast to the indirect limits from torsion balance equivalence principle (EP) tests [11, 12].
ULDM is well-motivated. It is consistent with the standard halo model (SHM) distribution of dark matter in a galaxy [12], and naturally occurs in BSM (beyond the SM) theories. The topologically complex vacua of string theory, for example, produces many ULDM candidates. These include moduli and dilatons [13, 14, 15, 16, 17, 18, 19] that couple to matter by making SM parameters depend on moduli fields. The moduli can acquire mass through supersymmetry (SUSY) breaking at $\sim$0.1 meV for TeV scale SUSY [13], or have much lighter masses due to loop factors and small coefficients (e.g. for the electron Yukawa modulus). The electron mass and charge ($m_{\rm e}$ and ($-e$), and the fine structure constant ($\alpha$), vary in time ($t$) and space ($\mathbf{x}$) in proportion to the ULDM field $\phi(t,\mathbf{x})$,
@@
Couplings to ULDM change $\alpha$ and $m_{\rm e}$, and hence the size of atoms and chemical bonds that go as the Bohr radius, $a_{{\scriptscriptstyle{\rm B}}}\equiv\hbar/(c\,m_{\rm e}\,\alpha)$. The effective ULDM “strain” driving the length $L_{\rm cav}(t)$ of a rigid optical cavity is the fractional change in $a_{{\scriptscriptstyle{\rm B}}}$ [22, 23]
@@
The bounds from this first demonstration of a broadband cavity length comparison method are limited by detector noise that is an order of magnitude above the photon shot noise. Additional laser noise filtering and vibration isolation, along with increased laser power to reduce the shot noise, should make it possible to profit from the cryogenic environment and improve these limits by 3 to 4 orders of magnitude. A longer science cavity and a non-rigid reference cavity [22] would extend the reach to lower ULDM Compton frequencies. The greatly improved broadband sensitivity could be used to search for ULDM from a boson star explosion in a “bosenova” near Earth [9] or from binary black hole and neutron star mergers [37]. Switching from our current broadband methods to resonant methods that use the quality factors of $10^{4}$ already observed could bring very large increases in sensitivity. More precise characterization of the mechanical resonance frequencies would allow a much deeper search in future experiments at these particular frequencies (Fig. 7) [38, 39]. Future measurements could even exploit the material properties of cryogenic sapphire cavities to realize quality factor values as high as $10^{8}$ [25, 40].
@@
In conclusion, a broadband direct ULDM detector sets new lower limits on the coupling of ultralight dark matter (ULDM) and ordinary matter. The new bounds are up to two orders of magnitude lower for both the standard halo model (SHM) and for a relaxion star bound to Earth. Routes to sensitivity increases of 3 to 4 orders of magnitude now seem feasible, with even greater sensitivities at the mechanical resonance frequencies of the optical cavities. Perpendicular cavities could provide gravitational wave sensitivity at frequencies above the 10 kHz limit of current large detectors.

### v2

# Demonstration that Differential Length Changes of Optical Cavities are a Sensitive Probe for Ultralight Dark Matter

@@ Abstract
Measurements of differential length oscillations of Fabry-Perot cavities provide a sensitive and promising approach to searching for scalar ultralight dark matter (ULDM). The initial demonstration sets direct lower bounds that are one to two orders of magnitude lower for two model ULDM distributions – a standard galactic halo, and a relaxion star bound to Earth – ranging over a decade of ULDM mass and Compton frequency. The demonstration suggests how a much higher sensitivity to a much larger ULDM mass range can be obtained.
@@
This Letter reports a new approach to discovering or obtaining bounds on a well-motivated alternative – scalar, parity-even, ultralight dark matter (ULDM) that behaves as a classical wave [7]. Fig. 1 represents a demonstration apparatus that establishes new bounds that are lower by one to two orders of magnitude, over a tenfold change in ULDM mass and Compton frequency. Also indicated are paths to a much higher sensitivity over a much larger mass range, and the feasibility of operating at cryogenic temperatures to reduce the effect of thermal noise as this becomes increasingly important. The measurement detects differing length changes of two cryogenic cavities at the ULDM Compton frequency $f_{\phi}\equiv m_{\phi}c^{2}/(2\pi\hbar)$, for a ULDM mass $m_{\phi}$, Planck constant $\hbar$, and speed of light $c$. Sensitivity to the local dark matter density produces strong and direct limits on boson stars [8, 9] composed of relaxions – a proposed solution to the hierarchy problem [10] – a contrast to the indirect limits from torsion balance equivalence principle (EP) tests [11, 12].
ULDM is well-motivated, being consistent with the standard halo model (SHM) distribution of dark matter in a galaxy [12], and naturally occurring in BSM (beyond the SM) theories. The topologically complex vacua of string theory, for example, produce many ULDM candidates. These include moduli and dilatons [13, 14, 15, 16, 17, 18, 19] that couple to matter by making SM parameters depend on moduli fields. The moduli can acquire mass through supersymmetry (SUSY) breaking at $\sim 0.1\,{\rm meV}$ for TeV scale SUSY [13], or have much lighter masses due to loop factors and small coefficients (e.g. for the electron Yukawa modulus). The electron mass and charge ($m_{\rm e}$ and ($-e$), and the fine structure constant ($\alpha$), vary in time ($t$) and space ($\mathbf{x}$) in proportion to the ULDM field $\phi(t,\mathbf{x})$,
@@
Couplings to ULDM change $\alpha$ and $m_{\rm e}$, and hence the size of atoms and chemical bonds that go as the Bohr radius, $a_{{\scriptscriptstyle{\rm B}}}\equiv\hbar/(c\,m_{\rm e}\,\alpha)$. The effective ULDM “strain” driving the length $L_{\rm cav}(t)$ of a rigid optical cavity is the fractional change in $a_{{\scriptscriptstyle{\rm B}}}$ [22, 23, 24, 25]
@@
The demonstration measurement, limited by detector noise that is ten times the photon shot noise, suggests the possibility of very large improvements. The sensitivity could be improved by 3 to 4 orders of magnitude using laser noise filtering and vibration isolation (in addition to what has been demonstrated), along with increased laser power to reduce the shot noise, whereupon the thermal noise would be reduced by the demonstrated cryogenic environment. The frequency and mass reach can be greatly extended using a longer science cavity and/or a non-rigid reference cavity, down to frequencies of 100 Hz [24]. The greatly improved broadband sensitivity could be used to search for ULDM from a boson star explosion in a “bosenova” near Earth [9] or from binary black hole and neutron star mergers [39].
Switching from our current broadband methods to resonant methods that use the quality factors of $10^{4}$ already observed could bring very large increases in sensitivity. More precise characterization of the mechanical resonance frequencies would allow a much deeper search in future experiments at these particular frequencies (Fig. 7) [40, 41]. Future measurements could even exploit the material properties of cryogenic sapphire cavities to realize quality factor values as high as $10^{8}$ [27, 42].
@@
In conclusion, a demonstration of a fresh approach to broadband direct ULDM detection sets new lower limits on the coupling of ultralight dark matter (ULDM) and ordinary matter. The new bounds are up to two orders of magnitude lower for both the standard halo model (SHM) and for a relaxion star bound to Earth. Routes to sensitivity increases of 3 to 4 orders of magnitude, for a much broader range of ULDM masses now seem feasible, with even greater sensitivities at the mechanical resonance frequencies of the optical cavities. Perpendicular cavities could provide gravitational wave sensitivity at frequencies above the 10 kHz limit of current large detectors.